OpenStudy (anonymous):

Please help. :) Line t passes through the points (1, 4) and (4, 9). Line m passes through the point (3, 2) and is parallel to line t . What is the equation for line m?

OpenStudy (zepp):

Can you first find the slope of the first line?

OpenStudy (anonymous):

I'm totally lost on this problem...

OpenStudy (zepp):

A cool guy made a tutorial on that, http://openstudy.com/users/zepp#/updates/4fa3474fe4b029e9dc34125e :D

OpenStudy (anonymous):

It won't load, please help me omg. :(

OpenStudy (zepp):

\(\huge m = \frac{y_{2}-y_{1}}{x_{2}-x_{1}} \) Formula for the slope

OpenStudy (anonymous):

Find slope of the first line. As it is parallel to \(m\), find the equation of the line by plugging in the points into \(y=mx+b\) or \(y-y_1=m(x-x_1)\)